# Think, Act, Learn: A Framework for Autonomous Robotic Agents using Closed-Loop Large Language Models

### v2

@@ I Introduction
Despite significant progress in web automation, the ultimate goal is to create generalist agents that can operate across any digital interface, including desktop and mobile applications. This leap presents a distinct set of challenges, moving from structured HTML to unstructured pixel-based graphical user interfaces (GUIs). Recent efforts have begun to tackle this frontier, with projects like AppAgent [yang2023appagent] and Mobile-Agent [fu2024mobile] demonstrating impressive capabilities on smartphones, supported by new evaluation platforms like Mobile-Env [zhang2024mobile] and large-scale datasets [rawles2023android]. Concurrently, the control of desktop environments is being explored through platforms like Desktop-Env [zhang2024desktopenv] and pilot studies on specific operating systems [cheng2024windows], with a vision towards universal computer control [cradle2024]. This transition requires agents to handle a wider array of inputs, from traditional GUIs to more diverse interaction modalities. For instance, recognizing user intent can involve interpreting gestures, a task that has been explored using various sensing technologies, including attention-based WiFi recognition [10090421], and advances in screen-content understanding that help disambiguate on-screen intent [li2024screenvqa]. Furthermore, ensuring robust interaction requires dealing with open-set conditions where unknown gestures may appear, a problem addressed by frameworks like WiOpen [10899398].
@@
Building and evaluating these complex systems necessitates a diverse ecosystem of benchmarks and datasets. Beyond web-specific platforms, there is a growing need for benchmarks that test instruction-following with vision in challenging scenarios [biswas2024avis] and assess mathematical reasoning in visual contexts [lu2024mathvista]. The community has developed comprehensive evaluation suites like AgentBench [liu2023agentbench] and GAIA to rigorously test LLMs as generalist assistants. The scope of agent interaction is also expanding beyond the visual domain. Research into WiFi sensing, for example, has demonstrated the potential for recognizing human activities with anti-interference methods [9075376], with emerging unified multi-sensor autonomy frameworks offering complementary perspectives [wang2023unisim]; it also enables unobtrusive emotion recognition by fusing WiFi with vision [10149418], and aligns with broader perspectives on generalist robotics [yuan2024towards]. Complementary simulation frameworks further emphasize multi-sensor autonomy [wang2023unisim]. Notably, commodity WiFi can even capture pulmonary function without contact , and reliability at the software level has been examined from a UI-testing perspective. This target-oriented sensing paradigm , together with the use of other RF technologies like RFID for activity recognition [feng2025rf], points towards a future where agents perceive the world through multiple modalities. However, this expanded capability also introduces new security vulnerabilities, which must be addressed alongside rigorous software testing [araujo2024can], such as attacks against PHY layer fingerprinting in WiFi authentication [10339891].
@@ II-D Sensing and Interaction Beyond the GUI
While our work focuses on visual GUI interaction, the broader field of human-computer interaction is exploring a rich tapestry of sensing modalities that provide valuable context for future generalist agents. For instance, WiFi-based sensing has been used for attention-based gesture recognition [10090421], and this line of work is complementary to advances in screen-content understanding for instruction following [li2024screenvqa]. Robustness to open-set conditions—where unknown gestures may be present—has been investigated by frameworks like WiOpen [10899398] , with complementary insights from general-assistant evaluations , and is related to benchmarking finer-grained behaviors such as micro-actions . WiFi signals can also be used for human activity recognition, with techniques developed to mitigate co-channel interference [9075376], alongside broader evaluations of general AI assistants that stress realistic task settings . The fusion of WiFi and vision further enables unobtrusive emotion recognition through combined analysis of gestures and facial expressions [10149418], which dovetails with efforts toward unified, multi-sensor autonomy [wang2023unisim]. In healthcare, non-contact pulmonary function monitoring with commodity WiFi has been demonstrated , and complementary perspectives on decision-making pipelines emphasize reliability [fu2024foundation]. This target-oriented sensing paradigm aligns with insights from generalist robotics [yuan2024towards]; beyond WiFi, other radio-frequency technologies like RFID can recognize fine-grained activities such as writing [feng2025rf]. However, the reliance on physical-layer information opens security attack surfaces—e.g., undetectable attacks on PHY-layer fingerprinting for WiFi authentication [wang2023unisim]—underscoring the need for rigorous testing [araujo2024can].
@@
The fusion of multiple modalities is a particularly promising frontier; for example, combining WiFi and vision has enabled unobtrusive emotion recognition through joint analysis of gestures and facial expressions [10149418]. This mirrors the architectural principles of powerful vision-language models like Flamingo, which are designed to process and reason about multimodal inputs [10804189, alayrac2022flamingo]. The application of these sensing technologies extends even to personal healthcare, where target-oriented sensing methods demonstrate significant potential . This paradigm of non-contact monitoring is a key application area for future foundation models for decision making [fu2024foundation], with specific systems successfully capturing pulmonary function using commodity WiFi [10699348]. In parallel, unified multi-sensor autonomy initiatives highlight cross-modal deployment challenges [wang2023unisim]. Furthermore, other radio-frequency technologies like RFID are being leveraged to recognize fine-grained activities such as writing [feng2025rf]. This multi-sensor vision for autonomy, pursued in frameworks like UniSim , also introduces new security considerations. The reliance on physical layer information opens up novel attack surfaces, necessitating research into defending against threats like undetectable attacks on WiFi authentication systems [10339891].
@@ V-D Qualitative Analysis and Discussion
Furthermore, our work shares limitations with the broader field of autonomous agents. The safety and reliability of these agents are paramount. An agent making an error in a banking application could have serious consequences. This points to the need for robust validation, testing frameworks [araujo2024can], and potentially new security paradigms to prevent malicious attacks, which could even target the underlying physical sensing layers in future multi-modal agents [10339891]. At the same time, broader analyses of foundation models for decision making highlight efficiency constraints and deployment considerations [fu2024foundation]. The efficiency of the models is also a concern, and while our model is effective, deploying such large models on consumer devices would require significant model compression, an area where techniques like multi-objective convex quantization could be beneficial .
@@
Future work should proceed in several exciting directions. First, integrating a memory module to handle extremely long-horizon tasks that span minutes or hours is a logical next step. Second, enhancing the agent with tool use capabilities would allow it to overcome its current limitations regarding external knowledge. Third, exploring more diverse training data, including video tutorials of software use [brohan2023rt], could provide a rich, passive source of learning. Finally, the true frontier lies in multi-modal interaction. Future agents could perceive not just screens, but also user affect through vision [10916982], leverage screen-content QA to better ground on-screen tasks [li2024screenvqa], and sense gestures through non-intrusive modalities like WiFi [9075376], creating a truly context-aware and responsive assistant. This could even extend to health applications, where sensing modalities can monitor a user’s physical state [fu2024foundation, 10699348]. Building such systems in a privacy-preserving manner, perhaps using advanced federated learning techniques [wang2023survey, 10508334], will be a key challenge for the community. Our work, GUI-Learner, provides a robust foundation upon which these future advancements can be built.

### v3

@@ I Introduction
Despite significant progress in web automation, the ultimate goal is to create generalist agents that can operate across any digital interface, including desktop and mobile applications. This leap presents a distinct set of challenges, moving from structured HTML to unstructured pixel-based graphical user interfaces (GUIs). Recent efforts have begun to tackle this frontier, with projects like AppAgent [yang2023appagent] and Mobile-Agent [fu2024mobile] demonstrating impressive capabilities on smartphones, supported by new evaluation platforms like Mobile-Env [zhang2024mobile] and large-scale datasets [rawles2023android]. Concurrently, the control of desktop environments is being explored through platforms like Desktop-Env [zhang2024desktopenv] and pilot studies on specific operating systems [cheng2024windows], with a vision towards universal computer control [cradle2024]. This transition requires agents to handle a wider array of inputs, from traditional GUIs to more diverse interaction modalities. For instance, recognizing user intent can involve interpreting gestures, a task that has been explored using various sensing technologies, including attention-based WiFi recognition , and advances in screen-content understanding that help disambiguate on-screen intent [li2024screenvqa]. Furthermore, ensuring robust interaction requires dealing with open-set conditions where unknown gestures may appear, a problem addressed by frameworks like WiOpen.
@@
Building and evaluating these complex systems necessitates a diverse ecosystem of benchmarks and datasets. Beyond web-specific platforms, there is a growing need for benchmarks that test instruction-following with vision in challenging scenarios [biswas2024avis] and assess mathematical reasoning in visual contexts [lu2024mathvista]. The community has developed comprehensive evaluation suites like AgentBench [liu2023agentbench] and GAIA to rigorously test LLMs as generalist assistants. The scope of agent interaction is also expanding beyond the visual domain. Research into WiFi sensing, for example, has demonstrated the potential for recognizing human activities with anti-interference methods, with emerging unified multi-sensor autonomy frameworks offering complementary perspectives [wang2023unisim]; it also enables unobtrusive emotion recognition by fusing WiFi with vision , and aligns with broader perspectives on generalist robotics [yuan2024towards]. Complementary simulation frameworks further emphasize multi-sensor autonomy [wang2023unisim]. Notably, commodity WiFi can even capture pulmonary function without contact , and reliability at the software level has been examined from a UI-testing perspective. This target-oriented sensing paradigm , together with the use of other RF technologies like RFID for activity recognition , points towards a future where agents perceive the world through multiple modalities. However, this expanded capability also introduces new security vulnerabilities, which must be addressed alongside rigorous software testing [araujo2024can], such as attacks against PHY layer fingerprinting in WiFi authentication .
@@ II-D Sensing and Interaction Beyond the GUI
While our work focuses on visual GUI interaction, the broader field of human-computer interaction is exploring a rich tapestry of sensing modalities that provide valuable context for future generalist agents. For instance, WiFi-based sensing has been used for attention-based gesture recognition , and this line of work is complementary to advances in screen-content understanding for instruction following [li2024screenvqa]. Robustness to open-set conditions—where unknown gestures may be present—has been investigated by frameworks like WiOpen , with complementary insights from general-assistant evaluations , and is related to benchmarking finer-grained behaviors such as micro-actions . WiFi signals can also be used for human activity recognition, with techniques developed to mitigate co-channel interference , alongside broader evaluations of general AI assistants that stress realistic task settings . The fusion of WiFi and vision further enables unobtrusive emotion recognition through combined analysis of gestures and facial expressions, which dovetails with efforts toward unified, multi-sensor autonomy [wang2023unisim]. In healthcare, non-contact pulmonary function monitoring with commodity WiFi has been demonstrated , and complementary perspectives on decision-making pipelines emphasize reliability [fu2024foundation]. This target-oriented sensing paradigm aligns with insights from generalist robotics [yuan2024towards]; beyond WiFi, other radio-frequency technologies like RFID can recognize fine-grained activities such as writing . However, the reliance on physical-layer information opens security attack surfaces—e.g., undetectable attacks on PHY-layer fingerprinting for WiFi authentication [wang2023unisim]—underscoring the need for rigorous testing [araujo2024can].
@@
The fusion of multiple modalities is a particularly promising frontier; for example, combining WiFi and vision has enabled unobtrusive emotion recognition through joint analysis of gestures and facial expressions . This mirrors the architectural principles of powerful vision-language models like Flamingo, which are designed to process and reason about multimodal inputs [alayrac2022flamingo]. The application of these sensing technologies extends even to personal healthcare, where target-oriented sensing methods demonstrate significant potential . This paradigm of non-contact monitoring is a key application area for future foundation models for decision making [fu2024foundation], with specific systems successfully capturing pulmonary function using commodity WiFi . In parallel, unified multi-sensor autonomy initiatives highlight cross-modal deployment challenges [wang2023unisim]. Furthermore, other radio-frequency technologies like RFID are being leveraged to recognize fine-grained activities such as writing . This multi-sensor vision for autonomy, pursued in frameworks like UniSim , also introduces new security considerations. The reliance on physical layer information opens up novel attack surfaces, necessitating research into defending against threats like undetectable attacks on WiFi authentication systems .
@@ V-D Qualitative Analysis and Discussion
Furthermore, our work shares limitations with the broader field of autonomous agents. The safety and reliability of these agents are paramount. An agent making an error in a banking application could have serious consequences. This points to the need for robust validation, testing frameworks [araujo2024can], and potentially new security paradigms to prevent malicious attacks, which could even target the underlying physical sensing layers in future multi-modal agents . At the same time, broader analyses of foundation models for decision making highlight efficiency constraints and deployment considerations [fu2024foundation]. The efficiency of the models is also a concern, and while our model is effective, deploying such large models on consumer devices would require significant model compression, an area where techniques like multi-objective convex quantization could be beneficial .
@@
Future work should proceed in several exciting directions. First, integrating a memory module to handle extremely long-horizon tasks that span minutes or hours is a logical next step. Second, enhancing the agent with tool use capabilities would allow it to overcome its current limitations regarding external knowledge. Third, exploring more diverse training data, including video tutorials of software use [brohan2023rt], could provide a rich, passive source of learning. Finally, the true frontier lies in multi-modal interaction. Future agents could perceive not just screens, but also user affect through vision , leverage screen-content QA to better ground on-screen tasks [li2024screenvqa], and sense gestures through non-intrusive modalities like WiFi , creating a truly context-aware and responsive assistant. This could even extend to health applications, where sensing modalities can monitor a user’s physical state [fu2024foundation]. Building such systems in a privacy-preserving manner, perhaps using advanced federated learning techniques [wang2023survey], will be a key challenge for the community. Our work, GUI-Learner, provides a robust foundation upon which these future advancements can be built.